### v2

# High-dimensional quantum key distribution using a multi-plane light converter

High-dimensional quantum key distribution using a multi-plane light converter
[…]
journal: opticajournalarticletype: Research Article
[…]
{abstract*}
High-dimensional quantum key distribution (QKD) offers higher information capacity and stronger resilience to noise compared to its binary counterpart. However, these advantages are often hindered by the difficulty of realizing the required high-dimensional measurements and transformations. Here, we implement a large-scale multi-plane light converter (MPLC) and program it as a high-dimensional mode sorter of spatial modes for QKD. Using the MPLC, we demonstrate five-dimensional QKD with six mutually unbiased bases and 25-dimensional QKD with two mutually unbiased bases in the same experimental setup. Furthermore, we propose a construction of pairs of mutually unbiased bases that are robust to experimental errors, with measurement complexity scaling only with the square root of the encoded dimension. This approach paves the way for QKD implementations in higher dimensions.
1 Introduction
Quantum key distribution (QKD) facilitates the sharing of a secret key between two parties, Alice and Bob, by encoding and sending quantum information between them[1, 2]. It is typically implemented by encoding quantum bits (qubits) onto the polarization or time-of-arrival of photons and was successfully demonstrated in a variety of experiments, including through long fiber-optic[3] or free-space[4, 5] links. Unfortunately, loss and noise in the optical link substantially reduce the achievable secure key rates in such systems[6], hindering their widespread use.
A promising approach for increasing the information capacity of QKD systems and improving their resilience to noise is through high-dimensional encoding[7, 8, 9]. In these systems, high-dimensional degrees of freedom are used to encode d-dimensional quantum states (qudits) that are then measured in two or more mutually unbiased bases (MUBs). High-dimensional quantum key distribution protocols have been demonstrated utilizing spatial [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], time-bin[26, 27, 28, 29, 30, 31], or time-energy [32, 33, 34, 35, 36, 37, 38, 39, 40] encoding. Spatial encoding is particularly attractive for high-dimensional QKD, as spatial light modulators (SLMs) allow efficient control over such high-dimensional states, which can then be transmitted through either multimode fibers[17, 41, 42], multi-core fibers[21, 23], or free-space links[43, 14].
However, the experimental complexity associated with measuring such spatially encoded qudits in multiple MUBs limits the performance of high-dimensional QKD. These measurements are often performed probabilistically: the state received by Bob is projected onto one out of $d$ possible states, making the measurement success rate inversely proportional to the dimension $d$. A better approach utilizes mode sorters, which direct each state to a different detector, ultimately enabling perfect measurements regardless of the dimension[12]. While mode sorters have been recently reported, realizing programmable mode sorters of spatial modes in multiple high-dimensional MUBs remains a major challenge.
Recently, multi-plane light converters (MPLCs) have emerged as a promising tool for realizing large, programmable mode sorters, using a cascade of optimized phase masks separated by free-space propagation[44]. The large number of programmable phase pixels in SLMs, together with the good connectivity provided by diffraction, make MPLCs a flexible platform that suits various applications[44, 45, 46, 47], ranging from classical communication[48] to quantum information processing[49, 50, 51, 52].
In this work, we use a record-large programmable MPLC with 10 phase planes to sort multiple MUBs of spatially entangled photons for high-dimensional QKD. We demonstrate two high-dimensional QKD schemes using the same experimental setup. We first implement a five-dimensional QKD scheme with sorted measurements in all six MUBs, obtaining a secure key rate of 1.57 bits per sifted photon through semidefinite programming on the tomographically-complete data[53]. We then perform 25-dimensional QKD in two MUBs. We take advantage of the flexibility and programmability of the MPLC to circumvent the fact that arbitrary mode sorters in a 25-dimensional space are beyond state-of-the-art experimental capabilities. Instead of the standard computational and discrete Fourier transform (DFT) bases, we use the MPLC to measure the photons in tailored MUBs with a complexity that scales only with the square root of the dimension. We further show that the structure of such MUBs leads to a non-uniform experimental error distribution, which is beneficial to the secure key rate, measured to be 0.8 bits per sifted photon in the experiment.
2 Mode sorting with a multi-plane light converter
Measurements in multiple MUBs are critical for QKD. Specifically, a mode sorter that can perform the required unitary transformations on the spatial modes that encode the state is crucial for increasing the secure key rate and must be realized for at least one MUB. This is in contrast to projective measurements with a success rate that is inversely proportional to the dimension.
We realize a high-dimensional programmable mode sorter in multiple MUBs using a 10-plane MPLC (fig.1a). For each MUB, the unitary transformation required for the change of basis is implemented by displaying different tailored phase masks on the MPLC planes. The phase masks are computed using the wavefront matching algorithm[48]. Examples of such phase masks used for 25-dimensional QKD are presented in fig.1b,c together with the summed intensity from all input modes at each plane of the MPLC.
In our experiment, we use the MPLC for high-dimensional QKD (fig.1a, see Supplementary). Alice produces pairs of spatially entangled photons via spontaneous parametric down-conversion (SPDC)[54]. The SPDC light passes at the far field through a binary amplitude mask with 50 apertures that define a 25-dimensional ’pixel-entangled’ quantum state[55]. To encode the information, Alice chooses a MUB by programming the appropriate unitary transformation between the set of input and output spatial modes using the MPLC. She then measures her photon at one of the output modes, which determines the state of the photon sent to Bob. To decode the information, Bob uses a second MPLC to sort the modes in a chosen (transposed) MUB and measures his photon at the output. For experimental convenience, we realize the two independent MPLCs by bouncing the entangled photons ten times between a single SLM and a mirror, where the top part of each phase mask is used by Alice and the bottom part by Bob.
Figure 1: Spatially-encoded QKD using multi-plane light conversion. (a) Pairs of spatially entangled photons are generated via SPDC and are transmitted through a binary amplitude mask with 50 apertures. Alice uses a programmable MPLC with ten phase masks to measure one of the photons in one of several MUBs. Its twin photon is transmitted to Bob, who uses his programmable MPLC to sort the different spatial modes in one of the MUBs. Measurements in which Alice and Bob choose the same basis contribute to the establishment of a secret key. Examples of such measurements in two MUBs of dimension 25 are presented in panels (b) and (c), where the optimized MPLC phase masks and the sum of intensities from all input modes are plotted for Alice (top half) and Bob (bottom half) for all planes.
3 Quantum key distribution in six mutually unbiased bases
To demonstrate the flexibility of MPLCs, we first consider five spatial modes per photon and perform five-dimensional QKD with measurements in all six MUBs[7, 56]. Along with the standard computational basis {|$\{\ket{n}\}_{n=1,\ldots,5}$ that can be measured without the MPLC, we choose five other MUBs {|$\left\{\ket{\psi^{(r)}_{k}}\right\}_{k=1,\ldots,5}$, with $k$ indexing the state and $r=2,\ldots,6$ the basis, such that[57]
|$\ket{\psi^{(r)}_{k}}=\frac{1}{\sqrt{5}}\sum^{5}_{n=1}e^{\frac{2\pi i}{5}[(k-1)(n-1)+(r-2)(n-1)^{2})]}\ket{n}$ (1)
For $r=2$, Eq. 1 is exactly the Discrete Fourier Transform (DFT) basis of dimension five. Note that the four other bases $r=3,\ldots,6$ only differ from the DFT basis by phases that are independent of the state $k$. Therefore, bases $r=3,\ldots,6$ can be mapped onto the DFT basis by applying mode-dependent phases to the spatial modes. For our pixel-entangled states, this is easily implemented using a single phase mask at the first plane of the MPLC. Measurements in these MUBs are thus realized by encoding the DFT transformation using the MPLC and superimposing mode-dependent phases at the first plane of the MPLC to choose the desired MUB (fig.2a).
The estimated probabilities for each of Bob’s measurement outcomes given a state sent by Alice, in all MUBs, are presented in fig.2b. From these probabilities, it is possible to compute a lower bound on the secure key rate. A rough lower bound can be obtained by applying the often-used simplifying assumption of identical depolarizing noise channels in all bases[7, 58, 16, 59]. This yields a secure key rate of $R=\log_{2}(d)-h_{d}(\frac{d+1}{d}E)-\frac{d+1}{d}E\log_{2}(d+1)=1.15\pm 0.05$ bits per sifted photon, where $h_{d}(x)=-x\log_{2}(\frac{x}{d-1})-(1-x)\log_{2}(1-x)$ is the high-dimensional Shannon entropy and $E=11\%$ is the arithmetic mean of the error rates in all bases.
However, a tighter estimation of the lower bound could be obtained through numerical techniques[60, 61, 31, 62, 53] using the full tomographically-complete data in fig.2b[53]. Indeed, using the method described in [53], we observe a gradual increase in the estimated key rate when taking into account more of the experimental data, both when Alice and Bob measure in the same basis and when measuring in different bases (see Supplementary information). We obtain a secure key rate of 1.57 for the full tomographically-complete data, which is significantly higher than the loose bound given by assuming depolarizing noise.
Figure 2: Five-dimensional QKD using a mode sorter in six MUBs. We program the MPLC to sort the photons in all six MUBs of dimension five. The bases are chosen such that one basis is the trivial computational basis, while the others are measured by adding different relative phases at the first plane of the MPLC, before realizing an optical discrete Fourier transform (DFT)(a). We measure the correlations between the photons in all bases and estimate the probability of each outcome in Bob’s measurement for each state sent by Alice (b). From these tomographically-complete measurements, we extract a secure key rate of 1.57 bits per sifted photon.
4 Experimentally-efficient 25-dimensional quantum key distribution
Next, we scale up to implement 25-dimensional QKD in two MUBs. As 25-dimensional arbitrary mode sorters are still out of reach[63], experimental demonstrations at this scale can use a simple mode sorter to measure in the computational basis, and perform projective measurements in a second MUB[64]. The fact that the first MUB is the computational one forces the second MUB to involve a fully connected transformation that interferes all d modes (e.g. d-by-d DFT), which makes its realization challenging.
However, thanks to the flexibility of the MPLC to sort modes in different bases, we can design tailored measurement bases that reduce the experimental complexity. Recently, a new approach to reduce the experimental complexity of four-dimensional QKD has been proposed, where instead of the computational and DFT bases, one uses two MUBs in which only pairs of modes are interfered[28, 21]. We generalize this approach to any dimension $d=p^{2m}$, where $p$ is a prime number and $m>0$ is an integer. In this case, we find two MUBs that involve the interference of only $\sqrt{d}$ modes, significantly reducing the complexity of the measurement (see Supplementary information).
We demonstrate this idea in dimension $d=5^{2}$ using our MPLC. The structure of the MUBs can be intuitively understood by placing the modes on a $\sqrt{d}$-by-$\sqrt{d}$ grid. Measurements in the two MUBs in this case are performed by applying $\sqrt{d}$-dimensional DFTs along the rows (MUB1, fig.3a) or columns (MUB2, fig.3b) of the grid. As any row and column only overlap at a single mode, these bases are mutually unbiased (see Supplementary information). We use the MPLC to experimentally realize such transformations on the 25 spatial modes of each photon and measure their correlations at the two MUBs (fig.3c).
Figure 3: 25-dimensional QKD in two MUBs. We encode a 25-dimensional quantum state in a grid of 25 spatially non-overlapping modes. Instead of using the standard computational and 25-dimensional DFT bases, we design two MUBs that are tailored to the experimental capabilities and only require the interference of sets of five spatial modes (a),(b). Measurements in these MUBs are performed by applying five-dimensional DFT transformations along the rows (a) or columns (b) of the 5x5 grid of modes using the MPLC. The experimental results are presented in (c), yielding a secure key rate of 0.8 bits per photon.
Interestingly, a non-uniform error distribution is clearly visible from the correlations in fig.3c. Along with uniform error sources such as accidental coincidence counts, the structure of the unitary transformation, that interferes blocks of $\sqrt{d}$ modes, leads to a block-biased error distribution. Non-uniform error distributions are also common in other mode sets such as Lagurre Gaussian modes in free space [65], linearly polarized modes in multimode fibers [66] or even time-bins [31]. In all cases, the block-biased noise behavior stems from the mode and MUB structures that dictate a lower probability of errors for distant modes that do not interfere. As the complexity of the semidefinite programming approach we used for $d=5$ prevents its implementation for $d=25$, in what follows, we utilize this block-biased noise structure to get a tighter analytical lower bound for the secure key rate.
In our experiment, the total error rate of $E_{t}=32.1\%$ is composed of a uniform error $E_{u}=7.3\%\approx 0.23E_{t}$ and a ’block’ error $E_{b}=24.8\%\approx 0.77E_{t}$. The total uniform error $E_{u}$ was estimated by summing the probabilities for measuring a state in one of the $d-\sqrt{d}$ states outside the ’block’ and multiplying by $\frac{d-1}{d-\sqrt{d}}$ to account for all $d-1$ error states. $E_{b}$ is then simply the remaining error probability within the $\sqrt{d}$-sized block.
We take this into account in the secure key rate lower bound and obtain:
$\begin{split}h_{d}(E_{u},E_{b})=-(1-E_{t})\log_{2}(1-E_{t})-\left(d-\sqrt{d}\right)\frac{E_{u}}{d-1}\log_{2}\left(\frac{E_{u}}{d-1}\right)\\ -\left(\sqrt{d}-1\right)\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)\log_{2}\left(\frac{E_{b}}{\sqrt{d}-1}+\frac{E_{u}}{d-1}\right)\end{split}$ (2)
In Eq. 2, the first term is for the measurement of the correct state, the second is for the $d-\sqrt{d}$ states outside the block, and the third is for the $\sqrt{d}-1$ possible error states within the block that are affected by both the uniform and block errors. Since uniform error maximizes the Shannon entropy, any non-uniformity in the error distribution increases the achievable secure key rate and total tolerable noise (fig.4a,b). The secure key rate is lower-bounded by $R=\log_{2}(d)-2h_{d}(E_{u},E_{b})$[67, 58, 59], yielding $0.8\pm 0.07$ secure bits per sifted photon in our experiment.
Figure 4: Secret key rates and error threshold. (a) The secret key rate per sifted photon in QKD is dependent on both the total error rate as well as the error distribution. In our 25-dimensional protocol, as a direct result of the block-like structure of the optical transformations, the errors are clearly not uniform and exhibit a strong block-like bias (orange curve). We therefore use a block-biased error model to calculate the secure key rate in this case, obtaining a rate of 0.8 bits per sifted photon (black marker). Non-uniform errors also increase the error threshold of the protocol (b). The error threshold increases significantly when the errors are block-biased, as in our experiment (orange), and even more so when they are completely within the block (green).
5 Conclusion
In this work, we have demonstrated high-dimensional QKD with spatially entangled photons by using MPLCs as programmable mode sorters. We have utilized the flexibility offered by our 10-plane MPLC to realize five-dimensional QKD with measurements in all six MUBs, and 25-dimensional QKD with measurements in two MUBs, in the same experimental setup.
For dimensions that are an even power of a prime number, we have further proposed pairs of MUBs for which the complexity of the experimental transformations only scales with the square root of the dimension. We have performed measurements in such MUBs for dimension 25 and shown that they lead to a non-uniform error distribution that allows for obtaining a tighter analytical lower bound for the secure key rate.
The main limitations of our proof-of-concept demonstration are the relatively high loss induced by the MPLC, which is currently in access of 10dB in our experiment (see Supplementary), and its slow refresh rate. These mainly result from the use of a programmable SLM and can be mitigated by using a constant phase mask instead[68]. While the full programmability of our MPLC will be lost in this case, our choice of MUBs ensures that simple switching between different MUBs could still be achieved, at a lower loss and a faster rate.
In the five-dimensional protocol, aside from the standard computational basis, switching between MUBs amounts to adding mode-dependent phases at a single plane, which can be realized using high-speed phase shifters. In the 25-dimensional protocol, the two MUBs only differ from one another by a rotation of the grid of modes by 90 degrees, which can be easily achieved using simple mirrors.
We thus believe that incorporating our demonstration of high-dimensional QKD using MPLCs into state-of-the-art spatial high-dimensional QKD systems [10, 11, 12, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 69, 24, 70] would pave the way towards high-rate long-range QKD.
[…]
Funding This research was funded by the Israeli Innovation Authority (Quantum Communication Consortium) and the Israel Science Foundation (grants No. 2497/21 and 2137/19). O.L. acknowledges the support of the Clore Scholars Programme of the Clore Israel Foundation. K.S. acknowledges the support of the Israeli Council for Higher Education. Y.B. and K.S. acknowledge the support of the Zuckerman STEM Leadership Program. The research of M.A. was supported by the European Union–Next Generation UE/MICIU/Plan de Recuperación, Transformación y Resiliencia/Junta de Castilla y León.
[…]
Disclosures
The authors declare no conflicts of interest.
[…]
Data availability Data underlying the results presented in this paper are available in Ref. [71].
[…]
Supplemental document See Supplement 1 for supporting content.